OpenStudy (anonymous):

WILL FAN AND MEDAL, Help please The surface area of a cube is 100 square centimeters. Estimate to the nearest tenth of a centimeter the side length of the cube using the formula S = 6s^2, where S represents surface area and s represents side length. Make sure to show all work and label your answer.

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

do you still need help?

OpenStudy (anonymous):

Yes @MelissaHolmes

OpenStudy (anonymous):

okay... here! :D so you want to solve for side length ....

OpenStudy (anonymous):

S = 6s^2 S is 100 so 100 = 6s^2

OpenStudy (anonymous):

100 = 6s^2 /6 /6 16.6 = s^2 sqrt(16.6) = s 4.07 = s -->your answer ...